# Machine-learning based high-bandwidth magnetic sensing

### v1

@@ Acknowledgments
Figure S1: LABEL:sub@fig:lfReal Normalised mean error of subsampled raster scans: blue circles show the error for 2 00020002\,0002 000 synthethic interpolated samples (the same of Fig. 2b in the main text). In purple squares, the same 2 00020002\,0002 000 synthetic samples that were subsampled but not interpolated. The red plot is similar to the blue, however all data samples are filtered to display eight non-overlapping Lorentzians. LABEL:sub@fig:nnSynth MAE of ML networks: in Purple squares, networks were trained on 10 0001000010\,00010 000 synthetic samples, and validated with 2 00020002\,0002 000 synthetic samples. In the latter, the networks were trained for specific number of data points (trained on subsampled datasets), i.e, each point on the graph depicts the result of a different network. The blue circles depict the result of one network, that was trained with 10 0001000010\,00010 000 synthetic samples, and validated with 2 00020002\,0002 000 synthetic samples, all at full length with 600 measurement points (the same of Fig. 2b in the main text). The subsampled test data was linearly interpolated to the full length (600 points) before using the network to get the predictions. The red plot is similar to the blue, however it was trained on a dataset in which all data samples displayed eight non-overlapping Lorentzians.
Fig. S1a reports the normalized error for the raster scans of different simulated datasets and at different subsampling levels. The blue dots are calcolated on 2 00020002\,0002 000 samples that were subsampled to the specified number of measurements and then linearly interpolated to the full size of 600 points. The red points are generated similarly with the exception that the samples are filtered to always have eight deeps without overlapping. In that case, the error is understandably lower. Finally, the purple squares are the normalized errors for 10 0001000010\,00010 000 not interpolated samples.
@@
Fig. S1b is equivalent to Fig. S1a but reports the MAE of the trained ML models. For the blue and red points the samples with less than 600 measurement points were linearly interpolated to the full size. Thus, in those cases we trained a single model with the full raster scan, and we tested it with the interpolated data. The purple squares represent the results of the models trained directly on the subsampled data. Thus for every point of the curve, a different model was trained and tested. The models used for the red points in Fig. S1b are trained with the same data used for the red points of Fig. S1a. Also in this case the error using the filtered data is understandably lower respect to the one that presents overlapping Lorentzians. However, it is remarkable that there is only a slightly difference between the two cases and that the ML models, contrarily to the raster scan, performs well also with the unfiltered data.

### v2

@@ Acknowledgments
Figure S1: LABEL:sub@fig:lfReal Normalised mean error of subsampled raster scans: blue circles show the error for 2 00020002\,0002 000 synthethic interpolated samples (the same of Fig. 3b in the main text). In red squares, the same 2 00020002\,0002 000 synthetic samples that were subsampled but not interpolated. The yellow plot with triangles is similar to the blue, however all data samples are filtered to display eight non-overlapping Lorentzians. LABEL:sub@fig:nnSynth MAE of ML networks: in red squares, networks were trained on 10 0001000010\,00010 000 synthetic samples, and validated with 2 00020002\,0002 000 synthetic samples. In the latter, the networks were trained for specific number of data points (trained on subsampled datasets), i.e, each point on the graph depicts the result of a different network. The blue circles depict the result of one network, that was trained with 10 0001000010\,00010 000 synthetic samples, and validated with 2 00020002\,0002 000 synthetic samples, all at full length with 600 measurement points (the same of Fig. 3b in the main text). The subsampled test data was linearly interpolated to the full length (600 points) before using the network to get the predictions. The yellow plot with triangles is similar to the blue, however it was trained on a dataset in which all data samples displayed eight non-overlapping Lorentzians.
@@
In Fig. S1a we report the normalized error for the raster scans of different simulated datasets and at different subsampling levels. The blue dots are calcolated on 2 00020002\,0002 000 samples that were subsampled to the specified number of measurements and then linearly interpolated to the full size of 600 points. The yellow triangles are generated similarly with the exception that the samples are filtered to always have eight resonances without overlapping. In that case, the error is understandably lower. Finally, the red squares are the normalized errors for 10 0001000010\,00010 000 not interpolated samples. Fig. S1b is equivalent to Fig. S1a but reports the MAE of the trained ML models. For the blue circles and yellow triangles the samples with less than 600 measurement points were linearly interpolated to the full size. Thus, in those cases we trained a single model with the full raster scan, and we tested it with the interpolated data. The red squares represent the results of the models trained directly on the subsampled data. Thus for every point of the curve, a different model was trained and tested. The models used for the yellow triangles in Fig. S1b are trained with the same data used for the yellow triangle of Fig. S1a. Also in this case the error using the filtered data is understandably lower respect to the one that presents overlapping Lorentzians. However, it is remarkable that there is only a slightly difference between the two cases and that the ML models, contrarily to the raster scan, performs well also with the unfiltered data.
Figure S2: Normalized error of raster scanning on samples with SNR values of 2.5 (blue circles), 4 (red squares), and 10 (yellow triangles).
In Fig. S2 we report the normalized error of raster scanning on the same samples used for the tests of Figs. 5a and 5b of the main paper. They are also the same values used in Fig. 2b where we report the not normalized error (the average of the values only where the raster scanning is able to find all the 8 resonances).
Finally, in Fig. S3, we present the evolution of the error on both the training set and a real test set for the models shown in Fig. 4 of the main text. Fig. S3a shows the training progress of the model trained with 10 0001000010\,00010 000 synthetic samples, while Fig. S3b corresponds to the model trained with 100 000100000100\,000100 000 samples. In both cases, the training error (along with the validation error shown in Fig. 4) decreases to below $0.5$ MHz. However, the test error on real data remains above $12$ MHz, indicating significant overfitting to the synthetic data.
Figure S3: Learning curves for the models shown in Fig. 4 of the main text, trained with 10 0001000010\,00010 000 LABEL:sub@fig:realTest1 and 100 000100000100\,000100 000 LABEL:sub@fig:realTest2 synthetic samples. For both models, the plots show the evolution of the error on the synthetic training set and on the same real test set used in Fig. 3 of the manuscript.